OpenStudy (ayyookyndall):

HELP!!! I AM SOO CONFUSED!! I need help writing the sine, cosine, and tangent ratios for

OpenStudy (ayyookyndall):



OpenStudy (unklerhaukus):

\[\sin A=\frac{\text{opposite side}}{\text{hypotenuse}}\]\[\cos A=\frac{\text{adjacent side}}{\text{hypotenuse}}\]\[\tan A=\frac{\text{opposite side}}{\text{adjacent side}}\]

OpenStudy (unklerhaukus):



OpenStudy (ayyookyndall):

Sin A = 3 / 8.5 cos A = 8 / 8.5 tan A = 3 / 8

OpenStudy (unklerhaukus):

\[\checkmark\\\checkmark\\\checkmark\]

OpenStudy (ayyookyndall):

sin A = 0.35294118 cos A = 0.94117647 tan A = 0.375

OpenStudy (unklerhaukus):

i would leave them in fraction form, but simplify a little \[3/8.5 = 6/17\\ 8/8.5=16/17\\3/8=3/8\]

OpenStudy (ayyookyndall):

Okay, can you help me with two more problem?

OpenStudy (unklerhaukus):

probably

OpenStudy (ayyookyndall):

Okay, so why would a measurement not be consistent if the we use the Converse of the Pythagorean Theorem?

OpenStudy (unklerhaukus):

i thought it was a right angled triangle,

OpenStudy (ayyookyndall):

The drawling below is what I am using. Under the ramp do you see a triangle? This will help you understand what I am talking about.

OpenStudy (unklerhaukus):

well according to that diagram (assuming the ground is flat), it should be right angled

OpenStudy (unklerhaukus):

if we use pythag. hypo = sqrt(adj^2+opp^2) but this doesn't quite give 8.50

OpenStudy (ayyookyndall):

Right, so my teacher said to make the sides as best as you can. (How long they would be in real life.) So, I marked adjacent side 8 ft and the opposite side 3 ft. but when you do the Pythagorean theorem the hypothenuse is 8.5

OpenStudy (unklerhaukus):

with pythag, i get 8.544

OpenStudy (unklerhaukus):

maybe we should have used the Law of cosines instead

OpenStudy (ayyookyndall):

And that doesnt go with the Converse of the Pythagorean Theorem so I was wondering why the measurement are not consisent..

OpenStudy (ayyookyndall):

The question only asked for me to do The Pythagorean Theorem and Converse of the Pythagorean Theorem.